MOLECULAR BEAM OPTICAL STARK MEASURMENTS OF ScS

T C. STEIMLE, \underline{D. M. GOODRIDGE} and A. J. MARR, Dept. of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287-1604.

The molecular beam optical LIF spectrum of the B^{2}Σ- X^{2}Σ^{+} band system of scandium monosulfide, ScS, has been recorded. The branch features associated with the low rotational levels were also recorded in the presence of a static electric field. The spectra were analyzed to produce magnetic hypefine parameters and permanent electric dipole moments. Comparisons with the theoretical predictions of Bauschlicher and Langhoff^{a}, and experimentally determined values for ScO^{b}are made.

aC.W. Bauschlicher, Jr. and S.R. Langhoff, J. Chem. Phys. 85, 5936 (1986)

bJ. Shirley, C. Scurlock, and T.C. Steimle, J. Chem. Phys. 93, 1568 (1990)